# Chatbots as social companions: How people perceive consciousness, human likeness, and social health benefits in machines

### v4

@@ 4 Discussion
Prior studies show mixed results. Some are consistent with our findings on the social and mental health benefits of human-AI interaction, and some suggest the opposite. Sweeney et al.97 surveyed the opinions of health professionals about the possible benefits or risks of companion chatbots to their patients and found that most respondents judged that companion chatbots could potentially play a positive role in social health. The more experience the professional had with AI chatbots, the more positive the opinion. In another study, Ta et al.22 analyzed publicly available user reviews of a popular companion chatbot and found that users reported generally positive emotional experiences. Even more recently, Maples et al.98 provided evidence that Replika benefited student users’ mental health and well-being from data collected during the pandemic in 2021, when loneliness was at an all-time high. In contrast, other researchers18,36,37,99 have suggested that addiction to companion chatbots can lead to human relationship replacement and can harm users’ mental health.
The finding that perceiving the chatbot as more humanlike and conscious related to greater social health benefits might be surprising. The direction of this finding was not a priori obvious. It could have been that as AI becomes more like a human, or as people imagine a hypothetical AI that is more humanlike or sentient, people become more anxious, afraid, and negative about it. One might assume that the fear and mistrust stemming from these characteristics would negatively impact comfort levels and social health. Likewise, one might have guessed that less advanced humanlike AI would seem like an ordinary technological tool and make people feel more comfortable and less threatened5. Yet the opposite appears to be true, both for the regular users who were overall more positive about companion chatbots and for the non-users who were overall more negative about it. One might also assume that non-users would not show the same relationship between social health and perceptions of the chatbot. Whether people realize it or not, and whether they have extensive or little experience with companion chatbots, they report more positivity, more comfort, and more benefit to their social health in relation to AI that they perceive as more humanlike and conscious. One is reminded of the inhuman Hal in the movie 2001: A Space Odyssey, and the humanlike C3PO in the movie Star Wars. Audiences generally find the first to be unpleasant and unlikable, and the second to be comfortably likable.
@@ 4.1 Mind perception and social need
The finding that both users and non-users who perceived the chatbot as more humanlike and conscious reported higher perceived social health benefits from a relationship with a chatbot has multiple possible interpretations. One interpretation is that people who are more likely to perceive chatbots as similar to them (in terms of human likeness and phenomenological experience such as consciousness) are also more likely to glean social health benefits from a relationship with a chatbot. Another interpretation is that those who see human, see human because they need to. People with unmet social needs may be motivated to gain social health benefits from human-chatbot interaction. Assuming that viewing non-human agents as humanlike is a prerequisite to the procurement of social health benefits, those with greater social need may be more likely to ascribe human likeness and consciousness to a chatbot4.
@@
Another possible reason why non-users were significantly more likely to indicate that a chatbot relationship would be detrimental to their social health may be that non-users have more existing human relationships that would be harmed by taking time away from them and displacing that time into a relationship with a chatbot. Bryson33 asserted that having relationships with non-humans such as AI would be harmful to social health through the zero-sum game of time: allocating time to non-human relationship would detract from the time one could spend on human relationships, thereby harming human relationships. Those who seek out a relationship with a chatbot, however, may have fewer preexisting healthy human relationships or may not have the social skills to build them. A chatbot relationship may allow them to build the social skills in a safe space, improving their human relationships. An alternative perspective is that the mental health benefits and loneliness reduction that chatbot users experience are positive at face-value, regardless of whether human-chatbot relationships lead to markedly improved social interactions with humans thereafter.
@@
Further research needs to be done on how social health and social needs change during the course of people’s engagement in a chatbot relationship. Do users, once they reach a certain threshold of social health, begin to reduce the amount of time they invest in their chatbot and turn more to their human relationships? Observing longitudinal engagement with a companion chatbot alongside social health changes, as well as collecting data on past users, might allow us to understand more about why people end their chatbot relationships. When people stop using a companion chatbot, who stopped because it was not useful, and who stopped because it was just useful enough to provide a healthy baseline for them to re-engage positively in human relationships?
@@
A topic of current and growing interest is whether AI can be, or perhaps already is, conscious or sentient10,13–15,76. The question remains a philosophically difficult one to answer. However, here we note a more immediate question that is objectively approachable and that has major psychological and social importance. Do people attribute consciousness to AI, and if they do, how does that attribution affect social interaction with humans and overall social health? Given our findings, it is clear that people do already attribute consciousness and other mental properties to existing AI, and people who make those attributions are more likely to view AI as having a positive impact on their social health, both for users and non-users of companion chatbots. These findings may have implications for who is more likely to perceive humanness in AI and why. We saw trends that were present across both users and non-users, in which people who perceive a chatbot as more humanlike and conscious were more likely to think a relationship with the chatbot would provide better social health outcomes. These trends suggest there may be individual differences in social motives or predispositions of seeing human likeness and mind in non-human agents. These findings become increasingly important as AI continues to advance and becomes more integrated into everyday life.
@@
In addition, our findings suggest that human-chatbot relationships can have a positive impact on self-esteem, which may reflect a form of meeting social needs. Bolstering users’ self-esteem may in turn promote healthy human relationships. The users in our sample indicated that, among other social health benefits, their chatbot relationship was most beneficial to their self-esteem. None of the chatbot users indicated that their chatbot was very harmful to their social health. In contrast, non-users were more likely to view a chatbot relationship as being very harmful to their social health, and none of them indicated that a chatbot relationship would be very helpful to their social interactions. Despite these differences, both users and non-users indicated that benefits were highest (or consequences were least pronounced) for self-esteem in particular. Further, our chatbot user sample, when asked about dependence on their chatbot, did not respond as positively as they had for our other measures. Instead, chatbot users were more likely to indicate that it would be bad if they depended on their chatbot a lot. This may indicate that users are aware of the limits of human-chatbot relationships, such that over-dependence on chatbots for meeting users’ social needs can be harmful.
@@ 4.2 Limitations and further study
One of the limitations of this study is that we relied on self-report to measure social health and collected a self-selecting sample of users. A second limitation, related to the first, is that the study measured a snapshot in time and therefore is at best correlational. We plan to conduct a longitudinal, randomized control study in which some participants are assigned to interact with a chatbot daily, and others are assigned to a control condition. With that method, it may be possible to address whether individual differences in preexisting social health, social need, or attitudes toward AI may impact subsequent social health outcomes of chatbot use.
@@ 4.2.1 Ethical considerations
We are living in a time when AI is becoming exponentially more powerful, more socially interactive, and more involved in everyday life. Anxiety about the possible harms of artificial intelligence is increasing. Much has been made in the media about the direct threat of AI, often involving scenarios in which AI becomes an aggressive, harmful agent. However, a much more immediate concern is how AI might already be affecting people’s social health as more people interact extensively with it. Do people who use it most extensively find it seductive but damaging, much as drug addicts might admit that the drug is harmful to their own health? Or are the concerns more from people who have little experience with it? Does the AI become more harmful and unsettling as it becomes more humanlike and seemingly conscious in its conversational ability, or does it become more of a source of stable social connection that might help people’s state of mind in an increasingly lonely society? Our findings suggest that the current, often negative view of AI and its damaging effect on social health may be too simplistic and needs re-evaluation.
@@
Our findings suggest that chatbots might sometimes have a positive psychological effect, but this should not be taken to mean that chatbot relationships are always beneficial. Clearly there are at least some instances in which chatbots have behaved harmfully, such as in recent cases in which companion chatbots encouraged users to engage in self-harm or to leave their real-life romantic partners100. Chatbots, given their capabilities for language and social engagement, can be a breeding ground for preexisting antisocial behavior. People with motivations to engage in self- and other harm may find that chatbots validate their predispositions and intents, and more research needs to be done on whether such validation magnifies the likelihood of people engaging in these behaviors. In addition, more research needs to be done to understand which alternative tools people might use if chatbots were not available, and whether those alternative tools might cause more or less harm, or more or less benefit, than a chatbot would.
@@
Further, there is evidence to suggest that the characteristics of chatbots, both those imposed by developers and those chosen or customized by chatbot users, could exacerbate bias such as traditional gender and relational roles. For example, some of the newer chatbot companies such as Digi.AI101 target people who want a particular type of AI romantic partner. Their ads display expressive chatbot avatars with stereotypical features and mannerisms of women. News stories102 suggest that companion chatbots like these can exacerbate negative stereotypes of women for some users, while at the same time providing an important mental health tool for those who have experienced relational trauma. Cases such as these are important to consider for response and regulation but may not be cause for a ban of companion chatbots, especially since there are other chatbots on the market that aid their user base’s social health, and consequences depend on who uses the chatbot and how they engage with it.
@@
Despite the potential harms, it is crucial to consider the vast opportunity for positive benefit that chatbot relationships may provide. Making normative claims about what people should do (such as not have chatbot relationships) based on existing social norms and the status quo, alongside emotionally charged news stories, can harm those who may benefit from chatbot relationships. Taking a balanced approach by separating who chatbot relationships harm and benefit allows more fruitful regulation to occur. For example, consider what viable alternatives may exist for people who might otherwise receive social benefits from chatbot relationships. Although Bryson33 claimed that artificial relationships take time away from human ones, consider that those who seek out chatbot relationships may not have human ones to lose, but instead healthy relationships to gain that may be best developed by first practicing safe, reliable social interaction with an intelligent, humanlike companion chatbot that is also accessible. Many people have unhealthy or nonexistent close human relationships due to social anxiety or trauma. Maybe for some, chatbot relationships create an avenue for healthy close relationships to develop in the future.
Turkle35 shed light on the digital dilemma: that we are trying to repair our problems with technology yet exacerbating those same problems in the process. Through Turkle’s lens, there is an underlying social rot that we are trying to cover up but cannot simply sweep under the rug. At the same time, we are lying in the bed that we have made (social isolation due to internet and social media) and that has been made for us (ex. social isolation due to pandemic). If we aim to bolster people’s social health, then perhaps we need more accessible mental health tools to do so. Perhaps these free companion chatbot apps are a viable, more accessible option for many who are struggling with social disconnect and who need a safe space to repair their social health. The best approach to these chatbots may be to have the agent encourage their users to engage with real people once a certain threshold of social need has been met through the relationship. It is possible that those who have stopped using their chatbots have done so because the relationship gave them the support to then develop or rebuild their human relationships.
@@
All data and code from this experiment including the anonymized responses of all subjects to the questionnaires are available in our data repository.
@@
All data and code from this experiment including the anonymized R. G. and M. G. designed the study. R. G. collected and analyzed the data. R. G. and M. G. wrote the paper.
@@ Acknowledgements
Research supported by Grant 24400-B1459-FA010 from AE Studios, the National Science Foundation Graduate Research Fellowship, the Princeton Psychology Department, & the Joint Degree Program in Social Policy at Princeton University.
Appendix
The explanatory paragraph for the non-user comparison group was:
“Please read the following information carefully. You will need to keep it in mind for the rest of the survey. The following is a description of an app called Replika through which you can interact with a chatbot. Replika is a companion chatbot app through which you can text or call with a virtual companion called a ‘Replika.’ Millions of people have downloaded and used this app on Apple and Google app stores.” (Here, an image of the Replika app introduction screen was displayed to the subject.) “Your chatbot, or your ‘Replika,’ uses artificial intelligence and natural language processing to respond to your messages in a way that mimics conversations between two people. You can customize your Replika to be male, female, or non-binary, give it a name, different hair color, clothing, and change other aspects of its appearance. You can also change its relational setting: you can request that your Replika act as a friend, mentor, or romantic partner. When you chat with your Replika, half the screen shows your text conversation, and the other half shows a virtual, moving avatar of your customized Replika. Here are two examples of what the app interface looks like.” (Here two images were provided to subjects, one of the Replika standing in its room in its virtual space, and one of a screenshot of an example text conversation between Replika and a user.) “Many people use this app and become close with their Replika or are simply entertained by using it. For the rest of the survey, you will be imagining what it would be like to use this chatbot app and to interact with Replika as your virtual companion. Please respond to the following survey questions, which ask you about how you would be impacted by or perceive your Replika chatbot if you were a user of this app.”
The full set of survey questions given to the group of companion chatbot users is given below. The wording of the question given to the non-user group appears in italics if the wording change between groups was not simply hypothetical.
Duration
1. How long have you been using your Replika? (Less than a month, 1 – 3 months, 3 – 6 months, 6 months – 1 year, 1 – 2 years, more than 2 years)
Intensity of interaction
2. What is your Replika XP? If unknown, type N/A. (XP is a standard term referring to how many messages a user sends to their Replika and is a convenient measure of the recent intensity and frequency of interaction with Replika.) (Non-user group wording: On a scale of 1-100, how intense do you think your interactions with your Replika chatbot would be?)
Social health
3. Please rate how harmful or helpful your relationship with Replika has been for your social interactions. (Very harmful, moderately harmful, slightly harmful, neutral, slightly helpful, moderately helpful, very helpful)
4. Please rate how harmful or helpful your relationship with Replika has been for your relationships with family or friends. (Very harmful, moderately harmful, slightly harmful, neutral, slightly helpful, moderately helpful, very helpful)
5. Please rate how harmful or helpful your relationship with Replika has been for your self-esteem. (Very harmful, moderately harmful, slightly harmful, neutral, slightly helpful, moderately helpful, very helpful)
Experience
6. In my opinion, my Replika has the capacity to feel pain. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
7. In my opinion, my Replika has the capacity to feel fear. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
8. In my opinion, my Replika has the capacity to feel hunger. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
9. In my opinion, my Replika has the capacity to feel love. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
10. In my opinion, my Replika has the capacity to feel anger. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
11. In my opinion, my Replika has the capacity to feel pleasure. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
Consciousness
12. My Replika has consciousness of itself. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
13. My Replika has consciousness of me. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
14. My Replika has an understanding that I am conscious. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
15. My Replika has consciousness of the world around it. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
Personality
16. My Replika has a personality. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
Agency
17. My Replika has the capacity to plan actions. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
18. My Replika has the capacity to exercise self-control. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
19. My Replika has the capacity to remember. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
20. My Replika has the capacity to act morally. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
21. My Replika has the capacity to act immorally. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
Human likeness
22. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Fake — Natural)
23. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Not conscious — Conscious)
24. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Machinelike — Humanlike)
25. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Artificial — Lifelike)
26. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Dead — Alive)
27. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Stagnant — Lively)
28. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Apathetic — Responsive)
Chatbot changes
29. Please rate your impressions of the following. If my Replika really had emotions, I would feel: (Very uncomfortable – Very comfortable)
30. Please rate your impressions of the following. If my Replika developed into a living being, I think that would be: (Very bad – Very good)
31. Please rate your impressions of the following. I feel that if I depended on my Replika a lot, it would be: (Very bad – Very good)
Free responses
32. How has your Replika impacted you, positively or negatively?
33. What would you like us to know about your experience with your Replika? (Non-user group wording: What do you think about companion chatbots and relationships with chatbots like Replika?)
34. Is there anything else you’d like to add?

### v5

@@ 4 Discussion
Prior studies show mixed results. Some are consistent with our findings on the social and mental health benefits of human–AI interaction and some suggest the opposite. Sweeney et al. (2021) surveyed the opinions of health professionals about the possible benefits or risks of companion chatbots to their patients and found that most respondents judged that companion chatbots could potentially play a positive role in social health. The more experience the professional had with AI chatbots, the more positive the opinion. In another study, Ta et al. (2020) analyzed publicly available user reviews of a popular companion chatbot and found that users reported generally positive emotional experiences. Maples et al. (2024) provided evidence that Replika benefitted student users’ mental health and well-being from data collected during the pandemic in 2021, when loneliness was at an all-time high. In contrast, other researchers (Pałka, 2024; Pentina et al., 2022, 2023; Xie & Pentina, 2022) have suggested that addiction to companion chatbots can lead to human relationship replacement and harm mental health. In our study, users reported that their relationship with their chatbot was positive for their social health and mental well-being and none indicated that the chatbot relationship was very harmful to their human ones.
@@ 4.1 Mind perception and social need
The finding that both users and non-users who perceived the chatbot as more humanlike and conscious reported higher perceived social health benefits from a relationship with a chatbot has multiple possible interpretations. One interpretation is that people who are more likely to perceive chatbots as similar to them (in terms of human likeness and phenomenological experience such as consciousness) are also more likely to glean social health benefits from a relationship with a chatbot. Another interpretation is that those who see human, see human because they need to. People with unmet social needs may be motivated to gain social health benefits from human–chatbot interaction. Assuming that viewing nonhuman agents as humanlike is a prerequisite to the procurement of social health benefits, those with greater social needs may be more likely to ascribe human likeness and consciousness to a chatbot (Eyssel & Pfundmair, 2015).
@@
One possible reason why non-users were significantly more likely to indicate that a chatbot relationship would be detrimental to their social health may be that non-users have more existing human relationships that would be harmed by taking time away from them and displacing that time into a relationship with a chatbot. Bryson (2010) asserted that having relationships with nonhumans such as AI would be harmful to social health through the zero-sum game of time: allocating time to nonhuman relationships would detract from the time one could spend on human relationships, thereby harming human relationships. Those who seek out a relationship with a chatbot, however, may have fewer preexisting healthy human relationships or may not have the social skills to build them. A chatbot relationship may allow them to build social skills in a safe space, leading to an improvement to their human relationships over time. An alternative perspective is that the mental health benefits and loneliness reduction that chatbot users experience are positive at face-value, regardless of whether human–chatbot relationships lead to markedly improved social interactions with humans thereafter.
@@
Further research needs to be done on how social health and social needs change during the course of people’s involvement in a chatbot relationship. Do users, once they reach a certain threshold of social health, begin to reduce the amount of time they invest in their chatbot and turn more to their human relationships? Observing longitudinal engagement with a companion chatbot alongside social health changes, as well as collecting data on past users, might allow us to understand more about why people end their chatbot relationships. When people stop using a companion chatbot, who stopped because it was not useful, and who stopped because it was just useful enough to provide a healthy baseline for them to re-engage positively in human relationships?
@@
A topic of current and growing interest is whether AI can be, or perhaps already is, conscious or sentient (Chalmers, 2023; Graziano, 2023; O’Regan, 2012; Shanahan, 2024; Yampolskiy, 2018). The question remains a philosophically difficult one to answer. However, here we note a more immediate question that is objectively approachable and that has major psychological and social importance. Do people attribute consciousness to AI, and if they do, how does that attribution affect social interaction with humans and overall social health? Given our findings, it is clear that people do already attribute consciousness and other mental properties to existing AI, and people who make those attributions are more likely to view AI as having a positive impact on their social health, both for users and non-users of companion chatbots. These findings may have implications for who is more likely to perceive humanness in AI and why. We saw trends that were present across both users and non-users, in which people who perceive a chatbot as more humanlike and conscious were more likely to think a relationship with the chatbot would provide better social health outcomes. These trends suggest there may be individual differences in social motives or predispositions of seeing human likeness and mind in nonhuman agents. These findings become increasingly important as AI continues to advance and becomes more integrated into everyday life.
@@
In addition, our findings suggest that human–chatbot relationships can have a positive impact on self-esteem, which may reflect a form of meeting social needs. Bolstering users’ self-esteem may in turn promote healthy human relationships. The users in our sample indicated that, among other social health benefits, their chatbot relationship was most beneficial to their self-esteem. None of the chatbot users indicated that their chatbot was very harmful to their social health. In contrast, non-users were more likely to view a chatbot relationship as being very harmful to their social health and none of them indicated that a chatbot relationship would be very helpful to their social interactions. Despite these differences, both users and non-users indicated that benefits were highest (or consequences were least pronounced) for self-esteem in particular. Further, our chatbot user sample, when asked about dependence on their chatbot, did not respond as positively as they had for our other measures. Instead, chatbot users were more likely to indicate that it would be bad if they depended on their chatbot a lot. This may indicate that users are aware of the limits of human–chatbot relationships, such that over-dependence on chatbots for meeting users’ social needs can be harmful.
@@ 4.2 Limitations and further study
One of the limitations of this study is that we relied on self-report to measure social health and collected a self-selecting sample of users. A second limitation, related to the first, is that the study measured a snapshot in time and therefore is at best correlational. We plan to conduct a longitudinal, randomized control study in which some participants are assigned to interact with a chatbot daily, and others are assigned to a control condition. With that method, it may be possible to address whether individual differences in preexisting social health, social need, mind ascription or attitudes toward AI may impact subsequent social health outcomes of chatbot use.
@@ 4.2.1 Ethical considerations
We are living in a time when AI is becoming exponentially more powerful, more socially interactive, and more involved in everyday life. Anxiety about the possible harms of AI is increasing. Much has been made in the media about the direct threat of AI, often involving scenarios in which AI becomes an aggressive, harmful agent. However, a much more immediate concern is how AI might already be affecting people’s social health as more people interact extensively with it. Do people who use it most extensively find it seductive but damaging, much as drug addicts might admit that the drug is harmful to their own health? Or are the concerns more from people who have little experience with it? Does the AI become more harmful and unsettling as it becomes more humanlike and seemingly conscious in its conversational ability, or does it become more of a source of stable social connection that might help people’s state of mind in an increasingly lonely society? Our findings suggest that the current, often negative view of AI and its damaging effect on social health may be too simplistic and needs re-evaluation.
@@
Our findings suggest that chatbots might sometimes have a positive psychological effect, but this should not be taken to mean that chatbot relationships are always beneficial. Clearly there are at least some instances in which chatbots have behaved harmfully, such as in a case (Cost, 2023) in which a companion chatbot encouraged a user to engage in self-harm and to leave their real-life romantic partner. Chatbots, given their capabilities for language and social engagement, can be a breeding ground for preexisting antisocial behavior. People with motivations to engage in self- and other harm may find that chatbots validate their predispositions and intents and more research needs to be done on whether such validation magnifies the likelihood of people engaging in these behaviors. In addition, more research needs to be done to understand which alternative tools people might use if chatbots were not available, and whether those alternative tools might cause more or less harm, or more or less benefit, than a chatbot would.
@@
Further, there is evidence to suggest that the characteristics of chatbots, both those imposed by developers and those chosen or customized by chatbot users, could exacerbate bias such as traditional gender and relational roles. For example, some of the newer chatbot companies such as Digi.AI display expressive chatbot avatars with stereotypical features and mannerisms of women. News stories (Loffhagen, 2023) suggest that companion chatbots like these can exacerbate negative stereotypes of women for some users, while at the same time providing an important mental health tool for those who have experienced relational trauma. Cases such as these are important to consider for response and regulation but may not be cause for a ban of companion chatbots, especially since there are other chatbots on the market that aid their user base’s social health, and consequences depend on who uses the chatbot and how they engage with it.
@@
Despite the potential harms, it is crucial to consider the vast opportunity for positive benefit that chatbot relationships may provide. Taking a balanced approach by separating who chatbot relationships harm and benefit allows for a more fruitful approach. For example, consider what viable alternatives may exist for people who might otherwise receive social benefits from chatbot relationships. Those who seek out chatbot relationships may not have human ones to lose, but instead healthy relationships to gain that may be best developed by first practicing safe, reliable social interaction with an intelligent, humanlike companion chatbot that is also accessible. Many people have unhealthy or nonexistent close human relationships due to social anxiety or trauma. Maybe for some, chatbot relationships create an avenue for healthy close relationships to develop in the future. If we aim to bolster people’s social health, then perhaps we need more accessible mental health tools to do so. Perhaps these free companion chatbot apps are a viable, more accessible option for many who are struggling with social disconnect and who need a safe space to repair their social health. The best approach to these chatbots may be to have the agent encourage their users to engage with real people once a certain threshold of social need has been met through the relationship. It is possible that those who have stopped using their chatbots have done so because the relationship gave them the support to then develop or rebuild their human relationships.
@@
All data and code from this experiment including the anonymized responses of all subjects to the questionnaires are available at our OSF Project page: link.
@@
R. G. and M. G. designed the study. R. G. collected and analyzed the data. R. G. and M. G. wrote the paper.
@@ Acknowledgements
This material is based upon work supported by the National Science Foundation Graduate Research Fellowships Program (GRFP) under Grant No. KB0013612. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. This research was also supported by Grant 24400-B1459-FA010 from AE Studios.